OpenStudy (fanduekisses):

Find the sum of the first 200 positive integers. a1=23 d=-2

OpenStudy (fanduekisses):

I have to find a200 right? a200=23 + (200-1)-2 right? is this correct?

OpenStudy (anonymous):

Key words, "Sum" and "Positive" terms.

OpenStudy (anonymous):

Sum meaning the addition of all terms found using the formula, positive terms only.

OpenStudy (anonymous):

Sum formula: \[S_n = \frac{ n(a_1-a_n) }{ 2 }\] Where Sn = the sum of the n terms. a1 = the first term an = the nth term

OpenStudy (anonymous):

You seem to have everything except for an which you know how to find.

OpenStudy (fanduekisses):

so \[S _{200}= 200(23-a _{n})/\]

OpenStudy (anonymous):

Correct.

OpenStudy (fanduekisses):

over 2

OpenStudy (fanduekisses):

I have to find an

OpenStudy (fanduekisses):

\[a _{200}\]

OpenStudy (anonymous):

You can use the formula for the Nth term:\[a_n = a_1+(n-1)d\]

OpenStudy (anonymous):

Correct

OpenStudy (fanduekisses):

\[a _{200}= 23 + (200-1)-2\]

jimthompson5910 (jim_thompson5910):

The formula should be \[\Large S_n = \frac{ n(a_{1}+a_{n}) }{ 2 }\] notice how It's "a1 + an" and not "a1 - an"

OpenStudy (fanduekisses):

Yes, but I need to find an first

OpenStudy (fanduekisses):

I get \[a _{200} = -375\]

jimthompson5910 (jim_thompson5910):

You want the sum of the first 200 terms so plug in n = 200 into the Sn sum formula \[\Large S_n = \frac{ n(a_{1}+a_{n}) }{ 2 }\] \[\Large S_{200} = \frac{ 200(a_{1}+a_{200}) }{ 2 }\] do you see where to go from here?

OpenStudy (fanduekisses):

is a200= -375 ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (fanduekisses):

I get -35200

jimthompson5910 (jim_thompson5910):

me too \[\Large S_{200} = -35200\]

OpenStudy (fanduekisses):

in the answer key it says it is 20,100

jimthompson5910 (jim_thompson5910):

The problem is "Find the sum of the first 200 positive integers. a1=23 d=-2" right?

jimthompson5910 (jim_thompson5910):

Are you sure d = -2?

OpenStudy (fanduekisses):

well, a1= 23 d=-2 is from the previous question, I assumed the question was referring to that

jimthompson5910 (jim_thompson5910):

oh so you want to add up 1, 2, 3, ..., 198, 199, 200 ???

jimthompson5910 (jim_thompson5910):

if so, then a1 = 1 a200 = 200 n = 200

OpenStudy (fanduekisses):

ohhh

OpenStudy (fanduekisses):

my fault D:

OpenStudy (fanduekisses):

thanks

jimthompson5910 (jim_thompson5910):

no worries and yw